OpenStudy (anonymous):

Could someone help me on this please, Rationalize the numerator = [ [squareroot3(a+b)+5] - [sqroot3a+5] ]/h

OpenStudy (anonymous):

is this square root or cubed root?

OpenStudy (anonymous):

\[\frac{\sqrt{3(a+b)+5}-\sqrt{3a+5}}{h}\]

OpenStudy (anonymous):

multiply top and bottom by the conjugate of the numerator, namely \[\sqrt{3(a+b)+5}+\sqrt{3a+5}\] the numerator will be \[3(a+b)+5-(3a+5)=3a+3b+5-3a-5=3b\]

OpenStudy (anonymous):

the denominator will be \[h(\sqrt{3(a+b)+5}+\sqrt{3a+5})\]

OpenStudy (anonymous):

if the "b" is supposed to be an 'h', which is what i suspect, you can cancel at this step otherwise you are done